Bob Hope was born 2 years after Clark Gable and 2 years before Henry Fonda. In 1951, after all three of them had their birthdays, the sum of their ages was 144. In what years were they born?
Answer by checkley75(3666)

X+(X+2)+(X+4)=144
3X+6=144
3X=144-6
3X=138
X=138/3
X=46
HENRY FORD BORN 1951-46=1905
BOB HOPE BORN 1951-48=1903
CLARK GABLE BORN 1951-50=1901